Suppose that the utility function for two products is given by U = x2y, and the budget constraint is 2x + 3y = 120. Find the values of x and y that maximize utility.


1
Expert's answer
2022-01-31T09:33:24-0500

The utility is maximized, when:

MUx/Px = MUy/Py and 2x + 3y = 120.

MUx = U'(x) = 2xy,

"MUy = U'(y) = x^2."

"2xy\/2 = x^2\/3,"

y = x/3,

2x + 3×x/3 = 120,

3x = 120,

x = 40 units,

y = 40/3 = 13 1/3 units.